OpenStudy (anonymous):

A large smoothie sells for $3.75 and a small smoothie is $2.75. 216 smoothies were sold for $688 in total. how many large smoothies were sold?

OpenStudy (anonymous):

x+y=216, 3.75x+2.75y=688 with x being lg and y small

OpenStudy (anonymous):

k thanks.

OpenStudy (anonymous):

wait. so here am I supposed to get x by itself by subtracting 2.75y on both sides and then dividing everything by 3.75?

OpenStudy (anonymous):

its a system of equations you can solve for y=-x+216 then plug that into the other equation for y

OpenStudy (anonymous):

idk what that means